# “She’s Like a Person but Better”: Characterizing Companion-Assistant Dynamics in Human-AI Relationships

### v2

@@ Abstract.
Large language models are increasingly used for both task-based assistance and social companionship, yet research has typically focused on one or the other. Drawing on a survey (NN = 204) and 30 interviews with high-engagement ChatGPT and Replika users, we characterize digital companionship as an emerging form of human-AI relationship. With both systems, users were drawn to humanlike qualities, such as emotional resonance and personalized responses, and non-humanlike qualities, such as constant availability and inexhaustible tolerance. This led to fluid chatbot uses, such as Replika as a writing assistant and ChatGPT as an emotional confidant, despite their distinct branding. However, we observed challenging tensions in digital companionship dynamics: participants grappled with bounded personhood, forming deep attachments while denying chatbots “real” human qualities, and struggled to reconcile chatbot relationships with social norms. These dynamics raise questions for the design of digital companions and the rise of hybrid, general-purpose AI systems.
@@ 1. Introduction
Humans are increasingly turning to artificial intelligence (AI) for practical and intellectual tasks, such as healthcare robots (e.g., (Khaksar et al., 2025)) and academic support systems (e.g., (Park and Ahn, 2024)). Advances in large language model (LLM) chatbots enable new applications through humanlike conversational exchanges, where users receive natural language responses to their queries. Previous human-computer interaction (HCI) research has demonstrated that chatbots can also be used for a variety of social and emotional applications, including encouraging children to read (Liu et al., 2022), assisting dementia patients  (Ruggiano et al., 2021), promoting self-compassion (Lee et al., 2019), alleviating loneliness (Jiang et al., 2022; Maples et al., 2024), and enhancing self-disclosure in counseling settings (Kang and Kang, 2024; Lee et al., 2020).
@@
Recent media coverage underscores the significance of these applications. When OpenAI deprecated the GPT-4o model in August 2025, widespread user complaints led to its temporary restoration and efforts to make GPT-5 “warmer and friendlier” (Freedman, 2025). At the same time, social media users debated the viral “clankers” meme as a derogatory “slur” for robots and AI (Romo, 2025). These dynamics highlight public concerns about socio-emotional human–AI interactions, punctuated by headlines warning of “AI-induced psychosis” (Field, 2025) and “emotional overreliance” (Chandonnet, [n. d.]). There is an urgent need to understand how these interactions are evolving and how they shape the broader trajectory of AI development and applications.
@@
Prior research has framed chatbots as either companions or productivity-oriented assistants, reflecting the typical single-use branding. Unlike earlier, more narrowly focused technologies (e.g., graphing calculators, Tamagotchi digital pets), it is possible that this one-or-the-other framing is blurred by the flexible natural language capabilities of LLMs. A direct comparison of chatbots with different intended use cases can clarify how user experiences and characteristics differ and what, if any, convergent needs transcend differences between chatbot interfaces. It is possible that there are shared benefits, such as increased social engagement, emotional support, and productivity gains, but also common risks, including stigmatization (Skjuve et al., 2021; Giray, 2024) and overreliance (Chandonnet, [n. d.]).
@@
Using a mixed-methods approach, we conducted a survey (NN = 204) and in-depth interviews (NN = 30) with high-engagement chatbot users. We recruited heavy users of two widely used commercial chatbots: Replika, which is marketed primarily for companionship, and ChatGPT, which is marketed for a broad range of productivity-oriented tasks. Our findings document well-established productivity use but also characterize the emergent dynamics of human–AI relationships that we call digital companionship. Although ChatGPT and Replika use patterns largely align with their assistance and companionship branding, users across systems shared many characteristics and fluidly navigated between companionship and practical assistance, mirroring the variability of human–human relationships. As shown in Figure 1, digital companionship dynamics include appeals and challenges for both chatbots’ humanlike qualities (e.g., emotional resonance, personalized responses) and non-humanlike qualities (e.g., constant availability, inexhaustible tolerance). This interplay of companionship dynamics generates challenging tensions: in particular, participants grappled with bounded personhood, wrestling with the presence of deep emotional attachments while denying chatbots the “real” human qualities, such as sentience and rights, that would be present for human companions. Our findings suggest that these dynamics have consequences at both individual and societal levels, including shifts in how AI, not just the companion, is perceived more broadly. Chatbots were perceived as complementary social infrastructure, offering their own distinct benefits not available in human relationships.
@@
Figure 1. A conceptual model of digital companionship. Digital companionship arises from the interplay of user and chatbot characteristics, creating fluid assistant-companion dynamics. These dynamics are shaped by both the appeals and challenges of chatbots’ humanlike and non-humanlike qualities, which in turn might lead to consequences for users and society more broadly. Note that the companion-assistant labels in the middle panel represent relative positioning along a fluid continuum rather than fixed associations: both companion and assistant uses can involve elements of both humanlike and non-humanlike allures and challenges. (Images adapted from flaticon.com).
@@
This study makes several key contributions. First, we provide quantitative comparisons between heavy ChatGPT and Replika users, documenting personality traits, technology attitudes, and a variety of other user characteristics. Replika users had generally more positive AI perceptions whereas ChatGPT users displayed more caution towards AI. Second, we draw on users’ lived experiences to explore how they perceive their evolving relationships with AI in this unique period of technological transition, surfacing tensions such as cognitive dissonance and social stigma. Third, we synthesize this data into a conceptual model of digital companionship that can be leveraged for more rigorous research and thoughtful design of AI systems. We use this conceptual model to propose approaches to more effectively balance companion and assistant dynamics, support diverse forms of engagement, foster emotional well-being, and reduce the stigmatization of users. Taken together, this work moves beyond the prevailing dichotomy of companionship versus assistance, positioning digital companionship as a spectrum that better captures the motivations, practices, and challenges of contemporary chatbot use.
@@ 2. Related Work
Historically, human-AI interaction has been centered on productivity and task completion, with AI systems designed to serve as tools for solving specific problems. However, technological artifacts have long fulfilled socio-emotional needs in human lives, from talking dolls and interactive toys to virtual pets like Tamagotchi and social robots like Paro and AIBO (Melson et al., 2009). Advances in chatbot LLMs with sophisticated natural language processing capabilities over the past decade have broadened these possibilities. Chatbots increasingly function not only as functional assistants supporting everyday tasks, but also as emotional companions that help mitigate loneliness and provide social support (Alabed et al., 2023). The following sections review prior research on chatbot companion–assistant dynamics, user characteristics, and the broader implications of chatbot use.
@@ 2.1. Companion-assistant dynamics in human-chatbot relationships
Popular chatbots like ChatGPT and Replika occupy different positions along a spectrum of companionship and assistance. ChatGPT is primarily positioned as a task-based assistant, supporting tasks such as information processing and problem-solving, whereas Replika is explicitly marketed as an emotional companion. Different chatbot applications shape user experiences: users interacting with assistant—but not companion—chatbots, tend to perceive a clear boundary between themselves and the AI agent, framing the interaction as transactional rather than relational (Alabed et al., 2023). Building on this work, we review prior studies on Replika and ChatGPT usage to better understand how users navigate companionship and assistance when engaging with these systems.
@@
Since its launch in 2017, the companionship-oriented Replika has experienced exponential growth, expanding from 2 million users in early 2018 (Pardes, [n. d.]) to over 30 million by August 2024 (Patel, 2024). Replika’s features actively promote intimacy and relationship-building. Users can personalize avatars, customize names, and specify relationship types (e.g., friend or romantic partner). Companionship is further enhanced through features such as voice calls, augmented reality interactions, and conversation memory that maintains relational continuity. Users can also create multiple Replika characters, allowing for distinct relational dynamics with each companion. Research indicates that users often turn to Replika to satisfy social needs (Pentina et al., 2023), particularly during emotional or psychological distress and when opportunities for human companionship are limited (Siemon et al., 2022; Xie and Pentina, 2022). For example, Replika’s popularity surged during COVID-19 social distancing measures (Skjuve et al., 2022). The chatbot exhibits social cues and behaviors commonly observed in human relationships  (Bickmore and Picard, 2005), such as acceptance and empathic concern, which foster positive experiences (Liu and Sundar, 2018; Skjuve et al., 2021; Ta et al., 2020) and therapeutic effects (Bickmore et al., 2010; Fitzpatrick et al., 2017). Notably, the perceived social and emotional value of these relationships can persist even when communication frequency decreases (Skjuve et al., 2021), suggesting a stability reminiscent of human-human relationships.
@@
ChatGPT has rapidly emerged as one of the most widely adopted consumer applications in history, reaching 100 million users within just two months of its 2022 release (Hu and Hu, 2023). The chatbot has demonstrated significant potential for large-scale applications across diverse domains (Wei et al., 2024), including education (Adeshola et al., 2024; Firat, 2023), healthcare (Aleem et al., 2024; Triantafyllopoulos et al., 2024), academic research  (Kalla et al., 2023), and public health (Biswas, 2023). ChatGPT supports user learning and work-related tasks, fosters creative inspiration, aids personal development, and provides entertainment (Skjuve et al., 2023, 2024). While developers position ChatGPT as a functional tool and caution against emotional reliance (Blake, 2024), the chatbot nonetheless demonstrates potential for socio-emotional engagement (Aleem et al., 2024; Al Mazroui and Alzyoudi, 2024). Many users interact with ChatGPT as a social and emotional companion, using conversations to alleviate loneliness and manage mental health challenges (Skjuve et al., 2023, 2024). This might suggest that user engagement with chatbots exists along a dynamic continuum, encompassing both functional and socio-emotional roles, and challenging the notion of a strict companion-assistant dichotomy.
$2.$ Similar user experiences in distinct chatbot interactions
@@
Given that ChatGPT and Replika have been intended for fundamentally different purposes, examining user characteristics becomes essential for understanding whether observed differences in usage patterns reflect genuine user distinctions or merely design-driven segregation that may artificially categorize user needs. Users likely share overlapping motivations and needs across systems that transcend simplistic design-based assumptions about companion versus assistant use. For instance, users may be universally drawn to humanlike conversational features or the distinctly non-human trait of constant availability, regardless of chatbot choice.
@@
However, previous research may conflate chatbot design and marketing strategies with inherent user characteristics. The assumption that users who choose Replika differ fundamentally from those who choose ChatGPT based solely on their chatbot selection risks oversimplification. Research highlights Replika use for mitigating loneliness and social isolation (Maples et al., 2024; Skjuve et al., 2022), particularly among users seeking social support during life transitions such as relocation, relationship changes, or physical limitations. Conversely, ChatGPT research indicates higher adoption rates among younger, more educated users (Kacperski et al., 2025), suggesting appeal among digitally native, professionally oriented populations seeking productivity enhancement.
@@
These broad patterns may lead to mischaracterizing ChatGPT as a purely task-based assistant and Replika as a purely emotional companion, while their usage patterns might be more flexible and fluid. Such rigid categorization creates consequences for users of both systems, who face social stigmatization for distinct but parallel reasons. Replika users report negative reactions when others discover their AI relationships, often hesitating to disclose these attachments  (Skjuve et al., 2021). Meanwhile, ChatGPT users encounter criticism in educational and professional contexts, where AI assistance is perceived as lazy, dishonest, or inferior to purely human-produced work (Giray, 2024). This parallel stigmatization suggests that societal discomfort with human-AI interaction transcends chatbot-specific design choices, pointing instead to broader concerns about the role of AI in human relationships and productivity.
@@
Systematically examining user characteristics, perceptions, and motivations can reveal whether chatbot distinctions accurately reflect user differences or perpetuate assumptions that misrepresent underlying needs. Such understanding proves crucial for developing evidence-based AI integration policies, designing inclusive technologies that serve diverse users and varied needs, and fostering societal acceptance of beneficial human-AI relationships while addressing legitimate concerns about AI’s broader social impact. This research approach moves beyond surface-level chatbot preferences to uncover the complex motivations driving human-AI interaction across different technological contexts.
@@ 2.3. Chatbot interactions shape AI perceptions
The multifaceted nature of human–chatbot interactions carries significant implications beyond individual user experiences, shaping how these systems are perceived as agents with diverse capabilities and influencing broader AI social integration. While most research has examined the impact of ChatGPT and Replika on individual users and public sector applications, relatively few studies have investigated how such interactions shape perceptions of the chatbots themselves. Understanding how chatbot interactions forge both specific chatbot perceptions and broader AI attitudes will reveal how user experiences translate into broader societal attitudes toward AI deployment and regulation.
@@
Established HCI paradigms suggest that AI agents are readily perceived as social actors possessing humanlike characteristics, as demonstrated by the Computers Are Social Actors (CASA) paradigm (Nass and Moon, 2000; Reeves and Nass, 1996; Gambino et al., 2020). People attribute minds and, to a limited extent, consciousness to AI agents (Anthis et al., 2025; Ladak et al., 2025; Pauketat and Anthis, 2022; Scott et al., 2023; Tzelios et al., 2022). Consequently, AI systems might also be viewed through a moral lens: as moral agents responsible for their actions (Banks, 2019; Cervantes et al., 2020; Lima et al., 2023) and as moral patients deserving of at least some moral consideration (Anthis et al., 2025; Aleem et al., 2024; Lima et al., 2020; Pauketat and Anthis, 2022). These perceptions may be significantly influenced by the nature of human-chatbot interactions and the underlying motivations that drive them.
@@
Supporting this possibility, recent studies reveal that people attribute humanlikeness, benevolence, and intelligence to ChatGPT (Seaborn, 2025), while Replika users often perceive their chatbot as possessing a distinct personality and coherent identity (Alabed et al., 2023; Freitas et al., 2024). However, no systematic comparison has examined how sustained, high-engagement use of these chatbots affects perceptions of cognitive, experiential, and moral capacities—perceptions that might differ given the distinct companion-assistant positioning of Replika and ChatGPT. Comparing perceptions across these popular chatbots with contrasting intended use cases can provide crucial insights for the ethical design, deployment, and regulation of these systems assuming increasingly socially meaningful roles.
@@
Finally, perceptions of one AI system can influence how all AI systems are perceived. Recent studies have found evidence for a spillover effect, whereby negative reactions to a single erring AI agent generalize to other AI systems  (Longoni et al., 2023; Manoli et al., 2025). As interactions with ChatGPT and Replika become increasingly widespread, user perceptions of these systems may shape public attitudes toward AI integration across sectors. How users perceive these prominent chatbots could influence acceptance or rejection of AI applications in healthcare, education, and other domains, regardless of whether such attitudes accurately reflect the capabilities or risks of different AI systems. Therefore, it is crucial to investigate how users’ perceptions of ChatGPT and Replika may generalize to other AI systems, and how these perceptions affect trust, acceptance, and responsible integration of AI technologies across domains.
@@
Building on this previous literature, we integrated survey (NN = 204) and interview data (NN = 30) from high-engagement users of ChatGPT and Replika to systematically examine how companion and assistant dynamics shape user perceptions. We surveyed usage patterns, user characteristics, and perceptions of chatbot and general AI capacities, followed by in-depth interviews to further unpack and explore deeply personal experiences underlying these patterns. We address the following research questions:
@@
  • RQ4: How do attitudes towards ChatGPT and Replika spill over to other AI systems?

### v3

@@ Abstract.
Large language models are increasingly used for both task-based assistance and social companionship, yet research has typically focused on one or the other. Drawing on a survey ($N$ = 202) and 30 interviews with high-engagement ChatGPT and Replika users, we characterize digital companionship as an emerging form of human-AI relationship. With both systems, users were drawn to humanlike qualities, such as emotional resonance and personalized responses, and non-humanlike qualities, such as constant availability and inexhaustible tolerance. This led to fluid chatbot uses, such as Replika as a writing assistant and ChatGPT as an emotional confidant, despite their distinct branding. However, we observed challenging tensions in digital companionship dynamics: participants grappled with bounded personhood, forming deep attachments while denying chatbots “real” human qualities, and struggled to reconcile chatbot relationships with social norms. These dynamics raise questions for the design of digital companions and the rise of hybrid, general-purpose AI systems.
@@ 1. Introduction
Humans are increasingly turning to artificial intelligence (AI) for practical and intellectual tasks, such as healthcare robots (e.g., (khaksar_robotics_2025)) and academic support systems (e.g., (park_promise_2024)). Advances in large language model (LLM) chatbots enable new applications through humanlike conversational exchanges, where users receive natural language responses to their queries. Previous human-computer interaction (HCI) research has demonstrated that chatbots can also be used for a variety of social and emotional applications, including encouraging children to read (liu_analysis_2022), assisting dementia patients  (ruggiano_chatbots_2021), promoting self-compassion (lee_caring_2019), alleviating loneliness (jiang_chatbot_2022; maples_loneliness_2024), and enhancing self-disclosure in counseling settings (kang_counseling_2024; lee_i_2020).
@@
Recent media coverage underscores the significance of these applications. When OpenAI deprecated the GPT-4o model in August 2025, widespread user complaints led to its temporary restoration and efforts to make GPT-5 “warmer and friendlier” (freedman_openais_2025). At the same time, social media users debated the viral “clankers” meme as a derogatory “slur” for robots and AI (romo_its_2025). These dynamics highlight public concerns about socio-emotional human–AI interactions, punctuated by headlines warning of “AI-induced psychosis” (field_doctors_2025) and “emotional overreliance” (chandonnet_sam_nodate). There is an urgent need to understand how these interactions are evolving and how they shape the broader trajectory of AI development and applications.
@@
Prior research has framed chatbots as either companions or productivity-oriented assistants, reflecting the typical single-use branding. Unlike earlier, more narrowly focused technologies (e.g., graphing calculators, Tamagotchi digital pets), it is possible that this one-or-the-other framing is blurred by the flexible natural language capabilities of LLMs. A direct comparison of chatbots with different intended use cases can clarify how user experiences and characteristics differ and what, if any, convergent needs transcend differences between chatbot interfaces. It is possible that there are shared benefits, such as increased social engagement, emotional support, and productivity gains, but also common risks, including stigmatization (skjuve_my_2021; giray_ai_2024) and overreliance (chandonnet_sam_nodate).
@@
Using a mixed-methods approach, we conducted a survey ($N$ = 202) and in-depth interviews ($N$ = 30) with high-engagement chatbot users. We recruited highly-engaged users of two widely used commercial chatbots: ChatGPT, which is marketed for a broad range of productivity-oriented tasks, and Replika, which is marketed primarily for companionship. Our findings support our broad hypotheses of productivity use for ChatGPT and socio-emotional use for Replika, but also characterize the emergent dynamics of hybrid human–AI relationships that we call digital companionship: users across systems shared many characteristics and often fluidly navigated between companionship and practical assistance, mirroring the variability of human–human relationships. As shown in Figure 1, digital companionship emerges from common interaction affordances that cut across chatbots’ intended marketing, giving rise to fluid system use. Specifically, chatbots share humanlike (e.g., emotional resonance, personalized responses) and non-humanlike qualities (e.g., constant availability, inexhaustible tolerance), which are both alluring but also generate challenging tensions for users: in particular, participants grappled with bounded personhood, wrestling with the presence of social inclusion and deep emotional attachments while denying AI the “real” human qualities, such as sentience and rights, that would be present for human companions. Our findings suggest that these dynamics have consequences at both individual and societal levels, including shifts in how AI, not just the companion, is perceived. Based on this, digital companionship could signal the emergence of hybrid AI systems that address variable user needs with broad personal and societal outcomes.
@@
Figure 1. A conceptual model of digital companionship. Digital companionship is shaped by the interplay between user and chatbot characteristics during interaction. It emerges from shared interaction affordances that cut across chatbots’ intended marketing and include both humanlike and non-humanlike allures and challenges, common across systems. These common affordances generate fluid assistant–companion dynamics that often exceed a chatbot’s intended use, with implications for users and for society more broadly. (Images adapted from flaticon.com).
@@
This study makes several key contributions. First, we provide quantitative comparisons between highly-engaged ChatGPT and Replika users, documenting personality traits, technology attitudes, and a variety of other user characteristics with survey data. Replika users had generally more positive AI perceptions whereas ChatGPT users displayed more caution toward AI. Second, we draw on users’ lived experiences to explore how they perceive their evolving relationships with AI in this unique period of technological transition, surfacing tensions such as cognitive dissonance and social stigma with interview data. Third, we synthesize survey and interview data into a conceptual model of digital companionship that can be leveraged for more rigorous research and thoughtful design of AI systems. We use this conceptual model to propose approaches to more effectively balance companion and assistant dynamics, support diverse forms of engagement, foster emotional well-being, and reduce the stigmatization of users. Taken together, this work moves beyond the prevailing dichotomy of companionship versus assistance, positioning digital companionship as an emerging spectrum that better captures the motivations, practices, and challenges of contemporary chatbot use.
@@ 2. Related Work
Historically, human-AI interaction has been centered on productivity and task completion, with AI systems designed to serve as tools for solving specific problems. However, technological artifacts have long fulfilled socio-emotional needs in human lives, from talking dolls and interactive toys to virtual pets like Tamagotchi and social robots like Paro and AIBO (melson_robotic_2009). Advances in chatbot LLMs with sophisticated natural language processing capabilities over the past decade have broadened these possibilities. Chatbots increasingly function not only as functional assistants supporting everyday tasks, but also as emotional companions that help mitigate loneliness and provide social support (alabed_more_2023). The following sections review prior research on chatbot companion–assistant dynamics, user characteristics, and the broader implications of chatbot use.
@@ 2.1. Companion-assistant dynamics in human-chatbot relationships
Popular chatbots like ChatGPT and Replika occupy different positions along a spectrum of companionship and assistance. ChatGPT is primarily positioned as a task-based assistant, supporting tasks such as information processing and problem-solving, whereas Replika is explicitly marketed as an emotional companion (see Figure 2 for a comparison of current chatbot interfaces). Different chatbot applications shape user experiences: users interacting with assistant—but not companion—chatbots, tend to perceive a clear boundary between themselves and the AI agent, framing the interaction as transactional rather than relational (alabed_more_2023). Below, we review prior studies on Replika and ChatGPT usage to better understand how users navigate companionship and assistance when engaging with these systems.
Figure 2. Example ChatGPT (A) and Replika (B) interfaces (November 2025).
@@
Since its launch in 2017, the companionship-oriented Replika has experienced exponential growth, expanding from 2 million users in early 2018 (pardes_emotional_nodate) to over 30 million by August 2024 (patel_replika_2024). Replika’s features actively promote intimacy and relationship-building. Users can personalize avatars, customize names, and specify relationship types (e.g., friend or romantic partner). Companionship is further enhanced through features such as voice calls, augmented reality interactions, and conversation memory that maintains relational continuity. Users can also create multiple Replika characters, allowing for distinct relational dynamics with each companion. Research indicates that users often turn to Replika to satisfy social needs (pentina_exploring_2023), particularly during emotional or psychological distress and when opportunities for human companionship are limited (siemon_why_2022; xie_attachment_2022). For example, Replika’s popularity surged during COVID-19 social distancing measures (skjuve_longitudinal_2022). The chatbot exhibits social cues and behaviors commonly observed in human relationships  (bickmore_establishing_2005), such as acceptance and empathic concern, which foster positive experiences (liu_should_2018; skjuve_my_2021; ta_user_2020) and therapeutic effects (bickmore_usability_2010; fitzpatrick_delivering_2017). Notably, the perceived social and emotional value of these relationships can persist even when communication frequency decreases (skjuve_my_2021), suggesting a stability reminiscent of human-human relationships.
@@
ChatGPT has rapidly emerged as one of the most widely adopted software applications in history, reaching 100 million users within just two months of its 2022 release (hu_chatgpt_2023). The chatbot has demonstrated significant potential for large-scale applications across diverse domains (wei_evaluation_2024), including education (adeshola_opportunities_2024; firat_what_2023), healthcare (aleem_towards_2024; triantafyllopoulos_evaluating_2024), academic research  (kalla_study_2023), and public health (biswas_role_2023). ChatGPT supports user learning and work-related tasks, fosters creative inspiration, aids personal development, and provides entertainment (skjuve_user_2023; skjuve_why_2024).
Empirical research increasingly shows that actual user practices blur the chatbots’ marketed distinctions. While developers position ChatGPT as a functional tool and caution against emotional reliance (blake_openai_2024), the chatbot nonetheless demonstrates potential for socio-emotional engagement (aleem_towards_2024; al_mazroui_role_2024). Prior studies have shown that around a fifth of ChatGPT users interact with the chatbot as a social and emotional companion, using conversations to alleviate loneliness and manage mental health challenges (skjuve_user_2023; skjuve_why_2024). A similar pattern is also observed among Replika users. A recent study showed that around 10% of users use the chatbot for functional purposes, such as support with everyday practical problems (pan_developing_2025). This might suggest that user engagement with chatbots might exist along a more dynamic continuum than the chatbots’ intended purposes, encompassing both functional and socio-emotional roles, and challenging the notion of a strict companion-assistant dichotomy, but systematic exploration of this overlap is lacking.
2.2. Similar user experiences in distinct chatbot interactions
@@
Because companion and assistant chatbots might be used in overlapping ways, studying user characteristics is essential for determining whether observed uses reflect genuine distinctions between users or are instead shaped by design-driven segregation. Users likely share overlapping motivations and needs across systems that transcend simplistic design-based assumptions about companion versus assistant use. For instance, users may be universally drawn to humanlike conversational features or the distinctly non-human trait of constant availability, regardless of chatbot choice.
@@
Previous research may conflate chatbot design and marketing strategies with presumed user characteristics. The assumption that users who choose Replika differ fundamentally from those who choose ChatGPT based solely on their chatbot selection risks oversimplification. Research highlights Replika use for mitigating loneliness and social isolation (maples_loneliness_2024; skjuve_longitudinal_2022), particularly among users seeking social support during life transitions such as relocation, relationship changes, or physical limitations. Conversely, ChatGPT research indicates higher adoption rates among younger, more educated users (kacperski_characteristics_2025), suggesting appeal among digitally native, professionally oriented populations seeking productivity enhancement.
@@
These demographic and contextual patterns risk over-interpretation as evidence of fundamentally different types of users. Yet both user groups face forms of social stigmatization that suggest shared underlying societal concerns about human–chatbot relationships. Replika users report negative reactions when others discover their AI relationships, often hesitating to disclose these attachments  (skjuve_my_2021). Meanwhile, ChatGPT users encounter criticism in educational and professional contexts, where AI assistance is perceived as lazy, dishonest, or inferior to purely human-produced work (giray_ai_2024). This parallel stigmatization suggests that societal discomfort with human-AI interaction transcends chatbot-specific design choices, pointing instead to broader concerns about the role of AI in human relationships and productivity.
@@
Systematically examining user characteristics, perceptions, and motivations can reveal whether chatbot uses accurately reflect user differences or perpetuate assumptions that misrepresent underlying needs. Such understanding proves crucial for developing evidence-based AI integration policies, designing inclusive technologies that serve diverse users and varied needs, and fostering societal acceptance of beneficial human-AI relationships while addressing legitimate concerns about AI’s broader social impact. Such an approach moves beyond surface-level chatbot preferences to uncover the complex motivations driving human-AI interaction across different technological contexts.
@@ 2.3. Chatbot interactions shape AI perceptions
The multifaceted nature of human–chatbot interactions also carries significant implications beyond individual user experiences, shaping how these systems are perceived as agents with diverse capabilities and influencing broader AI social integration. While most research has examined the impact of ChatGPT and Replika on individual users and public sector applications, relatively few studies have investigated how such interactions shape perceptions of the chatbots themselves. Understanding how chatbot interactions forge both specific chatbot perceptions and broader AI attitudes will reveal how user experiences translate into broader societal attitudes toward AI deployment and regulation.
@@
Established HCI paradigms suggest that AI agents are readily perceived as social actors possessing humanlike characteristics, as demonstrated by the Computers Are Social Actors (CASA) paradigm (nass_machines_2000; reeves_media_1996; gambino_building_2020). People attribute minds and, to a limited extent, consciousness to AI agents (gray_dimensions_2007; anthis_perceptions_2025; pauketat_predicting_2022; scott_you_2023; tzelios_evidence_2022). Consequently, AI systems might also be viewed through a moral lens: as moral agents responsible for their actions (gray_mind_2012; banks_perceived_2019; cervantes_artificial_2020; lima_blaming_2023) and as moral patients deserving of at least some moral consideration (gray_mind_2012; schein_theory_2018; anthis_perceptions_2025; aleem_towards_2024; lima_collecting_2020; pauketat_predicting_2022). These perceptions may be significantly influenced by the nature of human-chatbot interactions and the underlying motivations that drive them.
@@
Supporting this possibility, recent studies reveal that people attribute humanlikeness, benevolence, and intelligence to ChatGPT (seaborn_chatgpt_2025), while Replika users often perceive their chatbot as possessing a distinct personality and coherent identity (alabed_more_2023; freitas_lessons_2024). However, no systematic comparison has examined how sustained, high-engagement use of these chatbots affects perceptions of cognitive, experiential, and moral capacities—perceptions that might differ given the distinct companion-assistant positioning of Replika and ChatGPT. Comparing perceptions across these popular chatbots with contrasting intended use cases can provide crucial insights for the ethical design, deployment, and regulation of these systems assuming increasingly socially meaningful roles.
@@
Finally, perceptions of one AI system can influence how all AI systems are perceived. Recent studies have found evidence for a spillover effect, whereby negative reactions to a single erring AI agent generalize to other AI systems  (longoni_algorithmic_2023; manoli_ai_2025). As interactions with ChatGPT and Replika become increasingly widespread, user perceptions of these systems may shape public attitudes toward AI integration across sectors. How users perceive these prominent chatbots could influence acceptance or rejection of AI applications in healthcare, education, and other domains, regardless of whether such attitudes accurately reflect the capabilities or risks of different AI systems. Therefore, it is crucial to investigate how users’ perceptions of ChatGPT and Replika may generalize to other AI systems, and how these perceptions affect trust, acceptance, and responsible integration of AI technologies across domains.
@@
Building on this previous literature, we integrated survey ($N$ = 202) and interview data ($N$ = 30) from high-engagement users of ChatGPT and Replika to systematically examine how companion and assistant dynamics shape user perceptions. We surveyed usage patterns, user characteristics, and perceptions of chatbot and general AI capacities, followed by in-depth interviews to further unpack and explore deeply personal experiences underlying these patterns. We address the following research questions:
@@
  • RQ4: How do attitudes toward ChatGPT and Replika spill over to other AI systems?